# ObjectClear: Complete Object Removal via Object-Effect Attention

### v1

@@ C.4 More Comparisons on In-the-wild Data
Figure 15: Object Removal on OBER-Wild - Comparison between methods conditioned on Object-Effect Mask (others) and Object Mask (ours). Since existing methods struggle to remove object effects when provided with only the object mask, we supply them with our annotated object-effect masks for a fair comparison. Although these methods are able to remove shadows with the additional effect region, they often introduce undesirable changes to the original background. In contrast, ObjectClear effectively removes the object and its associated effects using only the object mask, while preserving the background content with high fidelity.
C.5 Limitations
While ObjectClear exhibits strong performance in removing objects and their associated effects, it still faces challenges in highly complex scenarios. Specifically, (1) when multiple shadows heavily overlap, the model struggles to accurately disentangle and localize the effect corresponding to the target object; and (2) in cases involving extremely large and long shadow regions, ObjectClear may fail to completely remove the effect, occasionally leaving behind residual artifacts. Enhancing the disentanglement of complex effect regions and improving the removal of large-scale effects remain important directions for future research.

### v2

@@ 8.8 More Comparisons on In-the-wild Data
Comparison with Image Inpainting Methods. We compare our method with state-of-the-art image inpainting methods, including PowerPaint [48], Attentive Eraser [34], DesignEdit [10], and RORem [18]. Since these inpainting methods often struggle to remove object-associated effects when only provided with an object mask, we supply them with our annotated Object-Effect Masks for fair comparison, while our ObjectClear uses only Object Masks as input. As shown in Fig. 20, although these inpainting methods can remove shadows with the guidance of additional effect regions, they frequently introduce undesirable changes to the original background. In contrast, ObjectClear effectively eliminates the target object and its associated effects using only the object mask, while preserving background content with high fidelity.
Comparison with Object Removal Methods. We further compare ObjectClear with object removal methods (OmniEraser [37] and OmniPaint [42]) under a unified input setting, all methods use Object Masks as input to ensure fairness. As illustrated in Fig. 19, our ObjectClear demonstrates superior stability compared to the two baselines: it not only successfully eliminates the target object and its associated effects but also avoids unintended changes to the original background, preserving background fidelity with high precision. To further highlight the background preservation advantage, we visualize the pixel-wise difference maps of the compared methods in Fig. 20. ObjectClear maintains minimal pixel differences in non-target background areas, which clearly validates its ability to preserve the original background while accurately removing the object and its effects.
Efficiency Comparison with Object Removal Methods. ObjectClear is significantly more efficient than both OmniEraser [37] and OmniPaint [42]. While the latter two are built on Flux ($\sim 12$B parameters), ObjectClear is based on SDXL ($\sim 3.7$B), making our model considerably lighter. As a result, ObjectClear runs over $5\times$ faster during inference, i.e., 1.63 s/image (ObjectClear) vs. 9.42 s/image (OmniPaint) at $512\times 512$ resolution on an A100 GPU.
8.9 Limitations
While ObjectClear exhibits strong performance in removing objects and their associated effects, it still faces challenges in highly complex scenarios. Specifically, in cases with overlapping shadows from multiple objects or complex lighting conditions, it can be difficult to disentangle which shadows belong to which objects. As a result, the model may fail to remove the shadows of the target object (Fig. 17a) or remove shadows of other objects (Fig. 17b). Effectively disentangling object-specific shadows in such complex scenes remains an important direction for future work.
Figure 17: Limitations. In complex scenes where multiple objects may produce overlapping or intertwined effects (e.g., shadows and reflections), our method can consistently remove the objects but may sometimes fail to precisely eliminate the associated effects. This results in either (a) under-removal of the target effect or (b) over-removal of effects belonging to nearby objects.
Figure 18: Comparison with Image Inpainting Models. Since some methods struggle to remove object effects when provided with only the object mask, we supply them with our annotated Object-Effect Mask for a fair comparison, while our ObjectClear uses only Object Mask as input. Although these methods are able to remove shadows with the additional effect region, they often introduce undesirable changes to the original background. In contrast, ObjectClear effectively removes the object and its associated effects using only the object mask, while preserving the background content with high fidelity.
Figure 19: Comparison with Object Removal Models. All compared methods use Object Masks as input for fair comparison. Our ObjectClear demonstrates superior stability: it not only successfully eliminates the target object and its associated effects, but also avoids unintended changes to the original background, preserving background fidelity with high precision.
Figure 20: Comparison of Object Removal Models via Difference Maps. We present results of representative object removal methods, and their pixel-wise difference maps relative to the input. ObjectClear successfully removes the target object and its associated shadows while preserving the background. This advantage is clearly reflected by the low pixel-wise differences in non-target areas.